OpenStudy (anonymous):

Find f(g) F(x)=x^2/3 g(x)=x^6

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

$$ F(g(x))=g(x)^{2/3} $$ Now substitute the following in for g(x): $$ \large g(x)=x^6 $$

OpenStudy (anonymous):

X^2/3

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

Is it x^4

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

yes!!

OpenStudy (anonymous):

So we will be multiplying the exponents

OpenStudy (anonymous):

What if it was g(f) same thing?

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

$$ \large (x^a)^b=x^{a\times b} $$ g(f) is not equal to f(g), generally unless they are inverses of each other But you follow the same process for calculations.